### v3

@@ 4.2 Asteroseismology of supernova progenitors: low-order p and g modes in Cep stars
The Cep class member with the richest set of multiplets so far is the M star HD 192575. This pulsator was discovered from TESS data covering the Northern Continuous Viewing Zone (N-CVZ) of the satellite. Its oscillation spectrum based on available TESS data covering about 3 years (with a 1-year gap in the middle) is shown in Fig. 6. This is an updated oscillation spectrum compared to the one in the discovery paper by Burssens et al. (2023) who used only the first 352 d of TESS data. Fig. 6 reveals a complete identification for the quintuplet indicated in blue, rather than an incomplete one in Burssens et al. (2023). This changes the avoided crossing with the other quintuplet indicated in orange in Fig. 6 and leads to a somewhat more evolved star (without affecting the other parameters from the forward modelling). Moreover, the left component of the quintuplet in panel b. indicated with “-2?” is now significant and may point to an asymmetry in that multiplet or to a frequency unrelated to the quintuplet. Rotation inversion of the securely identified multiplets of this star leads to a level of differentiality of about 1.5 between the core and envelope rotation (Vanlaer et al., in prep.), in line with the gradient shear layer solution found by Burssens et al. (2023) and the result from 2D stellar evolution modelling by Mombarg et al. (2023). The asymmetries in the splittings of the quintuplets of this star stand in sharp contrast to the symmetrical triplets. The detected asymmetries point to the presence of a core magnetic field in this supernova progenitor (Vanlaer et al., in prep.). This is of vast interest because it would deliver an asteroseismic calibration to 3D simulations of core-collapse supernova explosions as an important spin-off from asteroseismology of massive stars, in line with Fig. 1 already discussed in Sect. 1.
@@
Only few of the numerous modes active in the fastest rotators among the Cep stars have been unambiguously identified so far. Much remains to be done on that front, as is the case for the p modes in early-type Be pulsators (such as the CoRoT target HD 49330, Huat et al. 2009). This is not surprising, given the findings of Daszyńska-Daszkiewicz et al. (2002) who studied mode coupling among linear p modes in models of rapidly rotating Cep stars up to second order in . Their conclusion is essentially that mode identification is only possible if the inclination angle of the rotation axis is well measured – a situation hardly encountered in practice so far, although modern interferometry gives a promising perspective (Domiciano de Souza et al. 2014b; Bouchaud et al. 2020). Lack of mode identification is currently the major challenge to advance asteroseismology of fast rotating p-mode pulsators of spectral types O to B3, including the pulsating early-type OBe stars whose mode frequencies occur in a similar regime as to those of the Cep stars (Neiner et al. 2009). We anticipate major progress from asteroseismology of Cep stars and p-mode Be stars from applying 2nd- or 3rd-order perturbative methods in as in Suárez et al. (2009). Updated modelling of such an approach to all the stars with already identified modes, and even more so from future 2D or 3D stellar evolution models, is an opportunity we come back to at the end of this review.
@@ 4.3 Asteroseismic inferences on transport processes from modes of rotating Sct stars
Let us now move on to the case of slowly, moderately, and rapidly rotating Sct stars. An extensive catalogue of such pulsators was already available prior to the space era (Rodríguez & Breger 2001). With their masses between roughly 1.5 to 2.5 M, the majority of stars in this class have the most complex internal structures among dwarfs, with a fast rotating convective core and thin convective outer envelope requiring time-dependent convection theory in a non-adiabatic environment for a proper mode description (Dupret et al. 2005c, a; Antoci et al. 2014).
@@
Bowman (2016) gave an extensive summary of the observational properties of Sct stars as evaluated by the end of the nominal Kepler mission, highlighting that their oscillation spectra are characterised by strong amplitude modulation and moderate to fast rotation. This observational behaviour is readily understood in terms of the theory of non-linear mode coupling by Buchler et al. (1997) and Goupil et al. (1998). With the TESS mission ongoing, the class of Sct stars keeps increasing with thousands of members (Antoci et al. 2019) with well characterised global properties deduced from the Gaia space data (Murphy et al. 2019). Many of the Kepler Sct stars actually turned out to be hybrid Dor/Sct pulsators experiencing both high-order g modes and low-order p modes (Grigahcène et al. 2010; Uytterhoeven et al. 2011; Bowman et al. 2016; Li et al. 2020b). These are most interesting asteroseismic targets, as the combination of high-order g modes and low-order p modes offers probing power throughout the entire stellar interior, notably the rotational properties (Audenaert & Tkachenko 2022).
@@
The Ledoux approximation in Eq. (14) is still valid for a small fraction of the Sct pulsators. Slow rotation notably occurs among several of the Dor/Sct hybrid pulsators. The 1st-order Ledoux approximation was readily applied to the amplitude spectrum of the slowly rotating hybrid pulsator KIC 11145123 by Kurtz et al. (2014). This delivered the first Dor/Sct star with a determination of internal rotation. This star reveals nearly uniform rotation with a period of about 100 d and a slightly faster surface than core rotation. Models with atomic diffusion, including radiative levitation, offer an excellent explanation of the g modes and surface abundances for this hybrid pulsator (Mombarg et al. 2020). Nevertheless, it has been suggested that the star’s “weird” and slow rotation, along with its surface abundances deviating from solar values, may be due to a merger event (Takada-Hidai et al. 2017). While such a scenario would be no surprise following Fig. 2, another such slowly rotating hybrid Dor/Sct star, KIC 9244992, has solar-like surface abundances and very similar seismic properties. Its rotation near the surface deduced from p-mode splittings amounts to 66 d, which is slightly slower than the rotation period of 64 d in the near-core region measured from its g-mode splittings (Saio et al. 2015). Both pulsators have a mass of about 1.5 M and are at an advanced stage of main-sequence evolution, with a central hydrogen mass fraction of about 10%.
@@
Two additional similar cases of slowly rotating hybrid Dor/Sct pulsators were discovered in the 15.3 d eccentric orbit () binary KIC 10080943 by Keen et al. (2015); Schmid et al. (2015). Their orbital separation is wide enough to perform asteroseismic modelling as if the stars are single yet coeval, as done by Schmid & Aerts (2016). This delivered an age of 1.08 Gyr with a 1.67 M primary and a 1.60 M secondary having and 29%, respectively. Both stars have similar, small convective core overshooting and envelope mixing. The primary’s and secondary’s near-core rotation periods are 7.16 d and 11.05 d, respectively. This delivers spin parameters for the detected and identified dipole g modes between 0.20 and 0.26 for the primary, and between 0.14 and 0.24 for the secondary. These values place them well into the super-inertial regime (cf. Fig. 3). For both stars, the Coriolis force can still be treated perturbatively and the linear modes can be modelled from 1st-order Ledoux splitting (Schmid & Aerts 2016). The envelope rotation deduced from rotational splitting of p modes resulted in about 7.46 d and 8.20 d for the primary and secondary, respectively. This gives core-to-envelope levels of 0.96 and 1.35, respectively. Thus, the radial differential rotation of the secondary again reveals a slower core than envelope, just as for KIC 11145123. The centrifugal and tidal forces active in this binary result in only small perturbations, up to 1.7% for the former and below 0.04% for the latter. Thus, while the centrifugal force has a stronger influence on the oscillation modes than the tides, it is so small that a 1st-order treatment of the Coriolis force is justified from this perspective as well.
@@
The core-to-envelope rotation rates of these four Dor/Sct stars were found to be fully in line with angular momentum transport due to internal gravity waves, an interpretation reached from hydrodynamical simulations with appropriate levels of differentiality by Rogers (2015). However, these stars are not representative for the class of the Sct stars as a whole. Indeed, the large majority of these pulsators rotate so fast that even 2nd-order perturbation theories in are too limited according to evaluations of Eq. (15) for typical class members. For this reason, a more elaborate 3rd-order perturbation formalism taking into account the Coriolis force and the centrifugal flattening with its coupling with the Coriolis force, was developed by Soufi et al. (1998). For the full expressions of the operators for this 3rd-order perturbative theory in we refer to that paper and the corrections proposed by Karami (2008). So far this 3rd-order theory has hardly been applied to unravel the internal physics of fast rotating Sct stars. This is due to lack of proper mode identification for a sufficient number of modes, just as for the fast rotating Cep and early-type Be pulsators. Indeed, while Bedding et al. (2020) managed to overcome the hurdle of mode identification for a sample of 60 very young Sct stars from regular patterns delivering global parameters of these stars from their large frequency spacing, no rotational splitting was found for this sample. Pressure-mode frequency spacings alone cannot give tight constraints on transport processes to improve the input physics of the models.
@@
Most Sct stars do offer the potential to calibrate transport processes. However, this requires more theoretical work on non-linear mode coupling in order to interpret their complex frequency spectra, which include numerous combination frequencies in addition to a variety of frequency structures (Bowman et al. 2016). So far this complexity in the observations has prevented unambiguous mode identifications except partially for a few slow rotators (Zima et al. 2006; Lenz et al. 2008; Chen et al. 2017; Chen & Li 2017; Sun et al. 2023b) or overtone radial pulsators (Murphy et al. 2020; Daszyńska-Daszkiewicz et al. 2023). Already more than a decade ago, Breger et al. (2012) nicely illustrated that the low-frequency g modes and rotational signals are physically connected to the high-frequency p modes in the fast rotating hybrid Sct/Dor pulsator KIC 8054146. This Kepler target is a prototypical case of a non-linearly coupled mode pulsator according to Eq. (2) in the presence of fast rotation ( km s). Predictions for linear oscillation modes applied to non-rotating evolutionary models following the simplified Eq. (3) as computed by Murphy et al. (2023) are likely insufficient to identify the modes correctly. For hybrid Sct/Dor stars, non-linear asteroseismology by means of coupled mode equations following Eq. (2) as originally developed by Buchler et al. (1997) and Goupil et al. (1998) offers a way forward, provided that the theoretical formalisms used are generalised to include the Coriolis force to fit the low-frequency g modes as for SPB pulsators (discussed in the next Section, Lee 2012, 2022; Van Beeck et al. 2023).
@@
For the fast rotating pure p-mode pulsators among the Sct stars such as Altair (Bouchaud et al. 2020), the challenge is even greater, because the mode identification cannot rely on period spacing patterns as for the hybrid pulsators. Also, high-degree p modes are involved among the high-amplitude modes as Le Dizès et al. (2021) and Rieutord et al. (2023) have shown from combined MOST space photometry and high-resolution spectroscopy. High-degree mode asteroseismology at the level of fitting the frequencies of individually identified modes to calibrate transport processes has yet to be developed. For stars such as Altair, this must be done by using future 2D or 3D stellar evolution models such as those delivered by the ESTER code for stars of higher mass (Rieutord et al. 2016; Mombarg et al. 2023) coupled to 2D or 3D pulsation codes such as TOP (Reese et al. 2006, 2021) or ACOR (Ouazzani et al. 2012, 2015). Reese et al. (2009a) have shown that the acoustic p modes occurring in spherical stars are turned into island modes for rotationally deformed stars in 2D pulsation computations. As shown by Reese et al. (2009b), these island modes can potentially still be identified in observations, provided that another family of modes, namely chaotic modes, are absent in the frequency spectra. Moreover, Reese et al. (2021) have further shown from 2D adiabatic pulsation computations in the asymptotic regime of high radial order that island modes still show a so-called “pseudo” large frequency separation. The typical differences between actual numerically computed frequencies and those deduced from asymptotic approximations range from 5% to 12%. This relatively modest frequency difference should in principle safeguard the opportunity to identify island modes from pseudo large frequency separations predicted to occur in deformed p-mode pulsators. This is definitely a promising way forward to advance p-mode asteroseismology of rapidly rotating Sct, Cep, and p-mode Be pulsators. Currently these theoretical 2D pulsation predictions for p modes in deformed fast rotators have not yet been turned into practical applications.
@@ 4.4 Gravito-inertial asteroseismology: high-order g modes in moderate rotators of intermediate mass
Moving to the left in Fig. 3 we end up with dwarfs undergoing slow resonant internal gravity waves with periodicities of order days. Gravity-mode asteroseismology was kick-started from data assembled by the CoRoT space telescope (Auvergne et al. 2009), with the discovery of periodic deviations from a constant g-mode period spacing in the B3V star HD 50230 (Degroote et al. 2010). This pulsator’s g modes behave according to the theoretical predictions for non-rotating pulsators made before its discovery in the pedagogical paper by Miglio et al. (2008). Bringing this theory and the CoRoT data together allowed for an estimation of the mixing near its convective core. This SPB star is a prototype of the sub-class of slow rotators among g-mode pulsators of intermediate mass, including as well the SPB star KIC 10526294 whose amplitude spectrum was already shown in the bottom part of Fig. 4 and the Sct/Dor hybrids KIC 11145123 and KIC 9244992 discussed above. The core boundary mixing of such slowly rotating g-mode pulsators was meanwhile deduced in several studies (Wu et al. 2018; Wu & Li 2019; Wu et al. 2020; Mombarg et al. 2020; Pedersen et al. 2021).
@@
However, the majority of B and F g-mode pulsators are moderate to fast rotators whose high-order g mode periods are of similar order as their rotation period (cf. blue to orange stars indicated in Fig. 4). Most of the identified modes in the 63 well studied prototypical samples of Dor and SPB stars assembled by Aerts et al. (2021) indeed have frequencies in the co-rotating reference frame well below 2. Thus these modes are in the sub-inertial regime. In this case, it is no longer appropriate to treat the Coriolis force perturbatively in Eq. (2), not even for moderate rotators (cf. Fig. 3). On the other hand, the centrifugal force is less important than the Coriolis force to describe such waves, cf. Eq. (15). When the perturbation to the gravitational potential can be ignored (Cowling approximation), one can make the oscillation equations separable in by adopting the TAR for a uniformly rotating star (Lee & Saio 1987; Bildsten et al. 1996; Lee & Saio 1997; Townsend 2003b; Savonije 2005; Mathis 2009). In that case, the radial and azimuthal dependences are the same as for the non-rotating case, while the latitudinal () dependence is governed by Laplace’s tidal equations, whose solutions are written in terms of the Hough functions (Hough 1898) depending only on the spin parameter . Each eigensolution of the oscillation equations in the TAR is connected to the eigenvalue of the latitudinal equation, instead of as would occur in the component of the oscillation equations for the non-rotating case. The TAR is valid for modes with frequencies fulfilling , , and and whose is dominantly horizontal such that one can ignore the horizontal component of the rotation vector and assume . This is an excellent approximation for the majority of modes in the Dor and SPB stars. A modern public tool to compute oscillation modes of rotating spherical stars in the TAR is available in Townsend et al. (2018). Lee & Baraffe (1995), on the other hand, describe a method for mode calculations in the sub-inertial regime taking the full rotation vector into account, as well as the deformation of the star up to second order in .
@@
has been termed the buoyancy radius of the star (Miglio et al. 2008; Ouazzani et al. 2017). It characterises the size and other properties of the g-mode cavity. For dwarfs the g modes have their dominant probing power just outside the convective core at the deep bottom of the radiative envelope (see Aerts et al. 2021, for illustrative figures). Since has the dimension of a time and represents the characteristic period for the g modes of the star, it has also been termed the buoyancy travel time (Aerts 2021).
@@
The mode period spacings observed in an inertial frame, , can be computed from the mode periods expressed in an inertial frame () and in the corotating frame () following the transformation formula
@@
with the rotation period of the star. Equations (16) and (18) offer a powerful observable diagnostic tool to achieve the identification of detected modes from patterns occurring in the amplitude spectra. Such patterns allow us to estimate and via the involvement of the spin parameter in Eq. (16). This type of asteroseismic inference involving moderate or fast rotation does require long and uninterrupted data strings, in order to reach proper precision for the mode periods.
@@
Major efforts to assemble and interpret period spacing patterns for ensembles of gravito-inertial pulsators were undertaken from 4-yr Kepler light curves by Pápics et al. (2014), Pápics et al. (2015), Van Reeth et al. (2015b), Pápics et al. (2017), Szewczuk & Daszyńska-Daszkiewicz (2018), Li et al. (2019, 2020b), Szewczuk et al. (2021), and Pedersen et al. (2021). Being more than twice as long as the CoRoT data strings, the TESS light curves for stars in the CVZs also turn out to be suitable to hunt for such patterns (Garcia et al. 2022b, a). We anticipate the discovery of thousands more stars with such oscillation patterns in the future as the TESS light curves become longer. Overall, observed diagrams of versus mode periods in the inertial frame of an observer are already available for more than 700 g-mode pulsators, covering stars with masses in the range M and all rotation rates from almost zero to critical rotation.

### v4

@@ 4.2 Asteroseismology of supernova progenitors: Low-order p and g modes in Cep stars
The Cep class member with the richest set of multiplets so far is the M⊙ star HD 192575. This pulsator was discovered from TESS data covering the northern continuous viewing zone (N-CVZ) of the satellite. Its oscillation spectrum based on available TESS data covering about three years (with a one-year gap in the middle) is shown in Fig. 6. This is an updated oscillation spectrum compared to the one in the discovery paper by Burssens et al. (2023), who used only the first 352 d of TESS data. Figure 6 reveals a complete identification for the quintuplet (in blue), rather than an incomplete one in Burssens et al. (2023). This changes the avoided crossing with the other quintuplet indicated in Fig. 6 (in orange) and leads to a somewhat more evolved star (without affecting the other parameters from the forward modelling). Moreover, the left component of the quintuplet in panel b (labelled ‘-2?’) is now significant and may point to an asymmetry in that multiplet or to a frequency unrelated to the quintuplet. Rotation inversion of the securely identified multiplets of this star leads to a level of radial-differential rotation of about 1.5 between the core and envelope (Vanlaer et al. in prep.), in line with the -gradient shear layer solution found by Burssens et al. (2023) and the result from 2D stellar evolution modelling by Mombarg et al. (2023). The asymmetries in the splittings of the quintuplets of this star stand in sharp contrast to the symmetrical triplets. The detected asymmetries point to the presence of a core magnetic field in this supernova progenitor (Vanlaer et al. in prep.). This is of vast interest because it would deliver an asteroseismic calibration to 3D simulations of core-collapse supernova explosions as an important spin-off from the asteroseismology of massive stars, in line with Fig. 1 already discussed in Sect. 1.
Fritzewski et al. (2024c) brought the first ensemble asteroseismology for a population of 164 Cep stars discovered from Gaia DR3 by Gaia Collaboration et al. (2023a). The nature of these pulsators was confirmed from their light curves extracted from the TESS full frame images (FFIs) by Hey & Aerts (2024). We come back to their mode identification in Sect. 8.1, but point out here that this study increased the sample of Cep stars with a measurement of internal core-to-surface rotation substantially. The level of radial-differential rotation for 48 of these newly discovered pulsators ranges from 1 to 2, except for a few stars that have a value up to 4.
@@
Only a few of the numerous modes active in the fastest rotators among the Cep stars have been unambiguously identified so far. Much remains to be done on that front, as is the case for the p modes in early-type Be pulsators (such as the CoRoT target HD 49330, Huat et al. 2009). This is not surprising, given the findings of Daszyńska-Daszkiewicz et al. (2002) who studied mode coupling among linear p modes in models of rapidly rotating Cep stars up to second order in . Their conclusion is essentially that mode identification is only possible if the inclination angle of the rotation axis is well measured, a situation hardly encountered in practice, although modern interferometry gives a promising perspective (Domiciano de Souza et al. 2014; Bouchaud et al. 2020). Lack of mode identification is currently the major challenge to advance asteroseismology of fast-rotating p-mode pulsators of spectral types O to B3, including the pulsating early-type OBe stars whose mode frequencies occur in a similar regime as to those of the Cep stars (Neiner et al. 2009). We anticipate major progress from asteroseismology of Cep stars and p-mode Be stars from applying second- or third-order perturbative methods in as in Suárez et al. (2009). Updated modelling from such an approach to all the stars with already identified modes, and even more so from future 2D or 3D stellar evolution models, is an opportunity we come back to at the end of this review.
@@ 4.3 Asteroseismic inferences on transport processes from modes of rotating Sct stars
We now move on to the case of slowly, moderately, and rapidly rotating Sct stars. An extensive catalogue of these pulsators was already available prior to the space era (Rodríguez & Breger 2001). With their masses between roughly 1.5 and 2.5 M⊙, the majority of stars in this class have the most complex internal structures among dwarfs, with a fast-rotating convective core and thin convective outer envelope requiring time-dependent convection theory in a non-adiabatic environment for a proper mode description (Dupret et al. 2005c, a; Antoci et al. 2014).
@@
Bowman (2016) gave an extensive summary of the observational properties of Sct stars as evaluated by the end of the nominal Kepler mission, highlighting that their oscillation spectra are characterised by strong amplitude modulation and moderate to fast rotation. This observational behaviour is readily understood in terms of the theory of non-linear mode coupling by Buchler et al. (1997) and Goupil et al. (1998). With the TESS mission ongoing, the class of Sct stars keeps increasing with thousands of members (Antoci et al. 2019) with well characterised global properties deduced from the Gaia space data (Murphy et al. 2019). Many of the Kepler Sct stars actually turned out to be hybrid Dor–Sct pulsators experiencing both high-order g modes and low-order p modes (Grigahcène et al. 2010; Uytterhoeven et al. 2011; Bowman et al. 2016; Li et al. 2020b). These are extremely interesting asteroseismic targets as the combination of high-order g modes and low-order p modes offers probing power throughout the entire stellar interior, notably the rotational properties (Audenaert & Tkachenko 2022).
@@
The Ledoux approximation in Eq. (14) is still valid for a small fraction of the Sct pulsators. Slow rotation notably occurs among several of the Dor–Sct hybrid pulsators. The first-order Ledoux approximation was readily applied to the amplitude spectrum of the slow-rotating hybrid pulsator KIC 11145123 by Kurtz et al. (2014). This delivered the first Dor–Sct star with a determination of internal rotation. This star reveals nearly uniform rotation with a period of about 100 d and a slightly faster surface rotation compared to core rotation. Models with atomic diffusion, including radiative levitation, offer an excellent explanation of the g modes and surface abundances for this hybrid pulsator (Mombarg et al. 2020). Nevertheless, it has been suggested that the star’s ‘weird’ and slow rotation, along with its surface abundances deviating from solar values, may be due to a merger event (Takada-Hidai et al. 2017). While such a scenario should be no surprise (given Fig. 2), another such slowly rotating hybrid Dor–Sct star, KIC 9244992, has solar-like surface abundances and very similar seismic properties. Its rotation near the surface deduced from p-mode splittings amounts to 66 d, which is slightly slower than the rotation period of 64 d in the near-core region measured from its g-mode splittings (Saio et al. 2015). Both pulsators have a mass of about 1.5 M⊙ and are at an advanced stage of main-sequence evolution, with a central hydrogen mass fraction of about 10%.
@@
Two additional similar cases of slow-rotating hybrid Dor–Sct pulsators were discovered in the 15.3 d eccentric orbit () binary KIC 10080943 by Keen et al. (2015); Schmid et al. (2015). Their orbital separation is wide enough to perform asteroseismic modelling assuming the stars to be single yet coeval, as done by Schmid & Aerts (2016). This delivered an age of 1.08 Gyr with a 1.67 M⊙ primary and a 1.60 M⊙ secondary having and 29%, respectively. The two stars have similar small convective core overshooting and envelope mixing. The primary’s and secondary’s near-core rotation periods are 7.16 d and 11.05 d, respectively. This delivers spin parameters for the detected and identified dipole g modes between 0.20 and 0.26 for the primary, and between 0.14 and 0.24 for the secondary. These values place them well into the super-inertial regime (see Fig. 3). For both stars, the Coriolis force can still be treated perturbatively and the linear modes can be modelled from first-order Ledoux splitting (Schmid & Aerts 2016). The envelope rotation deduced from rotational splitting of p modes resulted in about 7.46 d and 8.20 d for the primary and secondary, respectively. This gives core-to-envelope levels of 0.96 and 1.35, respectively. Thus, the radial-differential rotation of the secondary again reveals a slower core than envelope, just as for KIC 11145123. The centrifugal and tidal forces active in this binary result in only small perturbations, up to 1.7% for the former and below 0.04% for the latter. Thus, the centrifugal force has a stronger influence on the oscillation modes than the tides.
@@
The core-to-envelope rotation rates of these four Dor–Sct stars were found to be fully in line with angular momentum transport due to internal gravity waves, an interpretation reached from hydrodynamical simulations with appropriate levels of core-to-envelope rotation by Rogers (2015). However, these stars are not representative of the Sct class as a whole. The large majority of these pulsators rotate so fast that even second-order perturbation theories in are too limited according to evaluations of Eq. (15) for typical class members. For this reason, a more elaborate third-order perturbation formalism taking into account the Coriolis force and the centrifugal flattening with its coupling with the Coriolis force, was developed by Soufi et al. (1998). For the full expressions of the operators for this third-order perturbative theory in we refer to that paper and the corrections proposed by Karami (2008). To date, this third-order theory has hardly been applied to unravel the internal physics of fast-rotating Sct stars. This is due to lack of proper mode identification for a sufficient number of modes, just as for the fast-rotating Cep and early-type Be pulsators. While Bedding et al. (2020) managed to overcome the hurdle of mode identification for a sample of 60 very young Sct stars from regular patterns delivering global parameters of these stars from their large frequency spacing, no rotational splitting was found for this sample. Pressure-mode frequency spacings alone cannot give tight constraints on transport processes to improve the input physics of the models.
@@
Most Sct stars do offer the potential to calibrate transport processes. However, this requires more theoretical work on non-linear mode coupling in order to interpret their complex frequency spectra, which include numerous combination frequencies in addition to a variety of frequency structures (Bowman et al. 2016). So far this complexity in the observations has prevented unambiguous mode identifications except partially for a few slow rotators (Zima et al. 2006; Lenz et al. 2008; Chen et al. 2017; Chen & Li 2017; Sun et al. 2023b) or overtone radial pulsators (Murphy et al. 2020; Daszyńska-Daszkiewicz et al. 2023). More than a decade ago, Breger et al. (2012) nicely illustrated that the low-frequency g modes and rotational signals are physically connected to the high-frequency p modes in the fast-rotating hybrid Sct–Dor pulsator KIC 8054146. This Kepler target is a prototypical case of a non-linearly coupled mode pulsator according to Eq. (2) in the presence of fast rotation ( km s-1). Predictions for linear oscillation modes applied to non-rotating evolutionary models following the simplified Eq. (3) as computed by Murphy et al. (2023) are likely insufficient to identify the modes correctly. For hybrid Sct–Dor stars, non-linear asteroseismology by means of coupled mode equations following Eq. (2), as originally developed by Buchler et al. (1997) and Goupil et al. (1998), offers a way forward, provided that the theoretical formalisms used are generalised to include the Coriolis force to fit the low-frequency g modes as for SPB pulsators (discussed in the next section; Lee 2012, 2022; Van Beeck et al. 2024).
@@
For the fast-rotating pure p-mode pulsators among the Sct stars (e.g. Altair (Bouchaud et al. 2020)), the challenge is even greater because the mode identification cannot rely on period spacing patterns as for the hybrid pulsators. High-degree p modes also occur among the high-amplitude modes, as Le Dizès et al. (2021) and Rieutord et al. (2023, 2024) have shown from combined MOST and TESS space photometry and high-resolution spectroscopy. High-degree mode asteroseismology at the level of fitting the frequencies of individually identified modes to calibrate transport processes has yet to be developed. For stars such as Altair, this must be done by using future 2D or 3D stellar evolution models, for example those delivered by the ESTER code for stars of higher mass (Rieutord et al. 2016; Mombarg et al. 2023) coupled to 2D or 3D pulsation codes such as TOP (Reese et al. 2006, 2021) or ACOR (Ouazzani et al. 2012, 2015).
Reese et al. (2009a) have shown that the acoustic p modes occurring in spherical stars are turned into island modes for rotationally deformed stars in 2D pulsation computations. As shown by Reese et al. (2009b), these island modes can potentially still be identified in observations, provided that another family of modes, namely chaotic modes, are absent in the frequency spectra. Moreover, Reese et al. (2021) have further shown from 2D adiabatic pulsation computations in the asymptotic regime of high radial order that island modes still show a pseudo large frequency separation. The typical differences between actual numerically computed frequencies and those deduced from asymptotic approximations range from 5% to 12%. This relatively modest frequency difference should in principle safeguard the opportunity to identify island modes from pseudo large frequency separations predicted to occur in deformed p-mode pulsators. This is definitely a promising way forward to advance p-mode asteroseismology of rapidly rotating Sct, Cep, and some of the Be pulsators. Currently these theoretical 2D pulsation predictions for p modes in deformed fast rotators have hardly been turned into practical applications, with the notable exception of Altair in Rieutord et al. (2024).
@@ 4.4 Gravito-inertial asteroseismology: High-order g modes in moderate rotators of intermediate mass
Moving to the left in Fig. 3 we end up with dwarfs undergoing slow resonant internal gravity waves with periodicities on the order of days. Gravity-mode asteroseismology was kick-started from data assembled by the CoRoT space telescope (Auvergne et al. 2009), with the discovery of periodic deviations from a constant g-mode period spacing in the B3V star HD 50230 (Degroote et al. 2010). This pulsator’s g modes behave according to the theoretical predictions for non-rotating pulsators made before its discovery in the pedagogical paper by Miglio et al. (2008). Bringing this theory and the CoRoT data together allowed an estimation of the mixing near its convective core. This SPB star is a prototype of the sub-class of slow rotators among g-mode pulsators of intermediate mass, including the SPB star KIC 10526294 whose amplitude spectrum is shown in the bottom part of Fig. 4 and the Sct–Dor hybrids KIC 11145123 and KIC 9244992 discussed above. The core boundary mixing of such slow-rotating g-mode pulsators have since been deduced in several studies (Wu et al. 2018; Wu & Li 2019; Wu et al. 2020; Mombarg et al. 2020; Pedersen et al. 2021).
@@
However, the majority of B and F g-mode pulsators are moderate to fast rotators whose high-order g-mode periods are of similar order to their rotation period (cf. blue to orange stars indicated in Fig. 4). Most of the identified modes in the 63 well-studied prototypical samples of Dor and SPB stars assembled by Aerts et al. (2021) indeed have frequencies in the co-rotating reference frame well below 2. Thus, these modes are in the sub-inertial regime. In this case it is no longer appropriate to treat the Coriolis force perturbatively in Eq. (2), not even for moderate rotators (see Fig. 3). On the other hand, the centrifugal force is less important than the Coriolis force to describe such waves (see Eq. (15)). When the perturbation to the gravitational potential can be ignored (the Cowling approximation), one can make the oscillation equations separable in by adopting the TAR for a uniformly rotating star (Lee & Saio 1987; Bildsten et al. 1996; Lee & Saio 1997; Townsend 2003b; Savonije 2005; Mathis 2009). In that case the radial and azimuthal dependences are the same as for the non-rotating case, while the latitudinal () dependence is governed by Laplace’s tidal equations, whose solutions are written in terms of the Hough functions (Hough 1898) depending only on the spin parameter . Each eigensolution of the oscillation equations in the TAR is connected to the eigenvalue of the latitudinal equation, instead of as would occur in the -component of the oscillation equations for the non-rotating case. The TAR is valid for modes with frequencies fulfilling , and , and whose is dominantly horizontal such that one can ignore the horizontal component of the rotation vector and assume . This is an excellent approximation for the majority of modes in the Dor and SPB stars. A modern public tool to compute oscillation modes of rotating spherical stars in the TAR is available in Townsend et al. (2018). Lee & Baraffe (1995), on the other hand, describe a method for mode calculations in the sub-inertial regime taking the full rotation vector into account, as well as the deformation of the star up to second order in .
@@
has been termed the buoyancy radius of the star (Miglio et al. 2008; Ouazzani et al. 2017). It characterises the size and other properties of the g-mode cavity. For dwarfs the g modes have their dominant probing power just outside the convective core at the deep bottom of the radiative envelope (see Aerts et al. 2021, for illustrations). Since has the dimension of a time and represents the characteristic period for the g modes of the star, it has also been termed the buoyancy travel time (Aerts 2021).
@@
The mode period spacings observed in an inertial frame () can be computed from the mode periods expressed in an inertial frame () and in the corotating frame () following the transformation formula
@@
with the rotation period of the star. Equations (16) and (18) offer a powerful observable diagnostic tool to achieve the identification of detected modes from patterns occurring in the amplitude spectra. Such patterns allow us to estimate and via the involvement of the spin parameter in Eq. (16). This type of asteroseismic inference involving moderate or fast rotation requires long and uninterrupted data strings in order to reach proper precision for the mode periods.
@@
Major efforts to assemble and interpret period spacing patterns for ensembles of gravito-inertial pulsators were undertaken from 4 yr Kepler light curves by Pápics et al. (2014), Pápics et al. (2015), Van Reeth et al. (2015b), Pápics et al. (2017), Szewczuk & Daszyńska-Daszkiewicz (2018), Li et al. (2019, 2020b), Szewczuk et al. (2021), and Pedersen et al. (2021). Being more than twice as long as the CoRoT data strings, the TESS light curves for stars in the CVZs also turn out to be suitable to hunt for such patterns (Garcia et al. 2022b, a). We anticipate the discovery of thousands more stars with such oscillation patterns in the future as the TESS light curves become longer. Overall, the observed diagrams of versus mode periods in the inertial frame of an observer are already available for more than 700 g-mode pulsators, covering stars with masses in the range M⊙ and all rotation rates from almost zero to critical rotation.